ii. Calculate LHV (lower heating value) ΔHc for octane and ethanol. Compare these values to those tabulated for methane, ethanol and octane on a mass basis. Q: What does Heat of Combustion indicate about fuels? Q: Compare the heat of combustion for the two fuels and comment about any significance relative to operation of an aircraft?
Answer #1

Lower heating value:-

For compounds containing carbon, hydrogen and oxygen (as many organic compounds do), a general combustion reaction equation will be:

CaHbOc + (a + ¼b - ½c) O2 → aCO2 (g) + ½bH2O (l) + heat of combustion

and the heat of combustion can be calculated from the standard heat of formation of all compounds taking part in the reaction:

ΔHc° = -aΔHf°(CO2,g) - ½b ΔHf°(H2O,l) + ΔHf°(CaHbOc) + (a + ¼b - ½c) ΔHf°(O2,g) = -a(- 393.51) - ½b(-285.830) + ΔHf°(CaHbOc) + (a + ¼b - ½c)*0

      = a(393.51) + b(142.915) + ΔHf°(CaHbOc)

For ethanol, the constants a, b and c are 2, 6 and 1, respectively, and the chemical equation for the combustion of ethanol:

C2H6O(l) + 3O2(g) → 2CO2(g)+ 3H2O(l)

The standard heat of formation of liquid ethanol, ΔHf°(C2H6O, l), is -277.6 kJ/mol.

The heat of combustion of ethanol, ΔHc°(C2H6O, l) = 2*393.51 + 6*142.915 + (-277.6) = 1366.91 kJ/mol. This can be converted to kJ per mass units:

The molweight of ethanol is (2*12.01 + 6*1.01 + 1*16.00) = 46.08 g/mol

The heat of combustion of ethanol, ΔHc°(C2H6O, l) = 1366.91[kJ/mol] *1000[g/kg] / 48.08 [g/mol] = 29664 kJ/kg ethanol = 29.7 MJ/kg

Similarly we can calculate for octane and it will be equal to 44.43 MJ/kg.

heat of combustion, heat released during combustion . In particular, it is the amount of heat released when a given amount (usually 1 mole ) of a combustible pure substance is burned to form incombustible products (e.g., water and carbon dioxide); this amount of heat is a characteristic of the substance. Heats of combustion are used as a basis for comparing the heating value of fuels, since the fuel that produces the greater amount of heat for a given cost is the more economic. Heats of combustion are also used in comparing the stabilities of chemical compounds. For example, if equal quantities of two isomeric hydrocarbons burn to produce equal amounts of carbon dioxide and water, the one releasing more energy (i.e., with the higher heat of combustion) is the less stable, since it was the more energetic in its compounded form.